# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v1

@@ 5.1.1 Missing Data
Handling missing data in machine learning, especially in the context of the PHM is significant, as incomplete data can significantly impact the model performance and prediction accuracy. The strategies to address missing data in PHM have evolved, encompassing a range of techniques from basic deletion to advanced imputation methods. This was evident in the 2021/2022 PHM EU competitions, where missing data posed significant challenges. In the fuse test bench dataset (2021 PHM EU), about 10% of the data was missing, and it was not evenly distributed across variables. Meanwhile, the SPI dataset(2022 PHM EU) had missing information in specific fields. To tackle these issues, various approaches were employed.
A straightforward strategy is the deletion of data points with missing values, such as listwise or pairwise deletion. For example, in the 2022 SPI dataset, instances having missing values in crucial identifiers like ”Panel_ID”, ”Figure_ID”, and ”Component_ID” were eliminated [68]. However, this approach can lead to the loss of valuable information to some degree. On the imputation front, several innovative methods were used in the 2021 PHM EU competition. One team applied the Last Observation Carried Forward (LOCF) method, coupled with backward filling, to address the gaps [12]. Interpolation was another technique used to estimate missing values, ensuring that cases with absent data did not skew the results. Additionally, Ramezani’s team introduced a novel approach, PLS-MV, a Partial Least Squares-based method for imputing missing values [55]. In addition to various imputation methods highlighted in the competitions, there remains scope for exploration in future research such as mean/median/multiple, KNN, regression imputation, maximum likelihood estimation, and machine learning-based approaches. These methods offer a promising avenue for further enhancing data quality and model reliability in PHM systems.
@@ 5.1.2 Data Imbalance
Data imbalance arises when the distribution of classes in a dataset is uneven. This issue is especially pronounced in PHM due to the rarity of failure events in comparison to data representing normal operations. Data imbalance can significantly undermine the performance of machine learning models, particularly in classification tasks. When trained on imbalanced data, models may become biased towards the majority class (normal operation) and may not effectively recognize the minority class (failure). This leads to poor performance in predicting failures, which may bring ineffective maintenance planning and unexpected downtimes. The fuse test bench dataset from the 2021 PHM EU competition and the printed circuit board dataset from the 2022 PHM EU competition exemplify this issue. The former displayed a stark imbalance with only 3-4 experiments representing each fault class against 70 experiments for the healthy class. Similarly, in the printed circuit board dataset, a staggering 95% of the SPI data was classified as healthy, highlighting a significant imbalance. The 2018 PHM NA competition also faced this challenge in the Ion Mill Etching (IME) dataset, where the faulty data is much less than normal operation data.
To counteract data imbalance in PHM, various strategies have been adopted. Resampling techniques are commonly used in competitions to adjust the dataset to balance the class distribution, either by oversampling the minority class or undersampling the majority class. Synthetic data generation is another approach, as demonstrated by a team in the 2021 PHM EU competition that used the Synthetic Minority Over-sampling Technique (SMOTE) to create synthetic samples of the minority class to balance the dataset [12]. Advanced algorithms also play a crucial role in addressing data imbalance. For instance, Tang’s team in the 2022 PHM EU competition used the Feeding Imbalance Ratio (FIR) to control the imbalance ratio in each mini-batch during neural network training [70]. Liu et al. employed transfer learning and domain adaptation techniques in the 2018 PHM NA competition to facilitate knowledge transfer across different fault modes, addressing the issue of insufficient data in specific faults [44]. Moreover, during the 2018 PHM NA competition, many teams utilized Random Forest, an ensemble learning method known for its proficiency in handling imbalanced data, by constructing a forest of decision trees. These approaches collectively demonstrate the diverse and innovative methods employed to mitigate the challenges posed by data imbalance in the PHM field.
@@ 5.1.3 Domain Shift
Domain shift refers to the changes in the data distribution between the training phase (source domain) and the real-world application phase (target domain) of the machine learning models. This phenomenon frequently occurs when models, initially trained on data from a specific set of machines or under certain conditions, are subsequently applied to different machines or varied operating conditions. Such a shift can markedly affect the performance and reliability of machine learning models in PHM, making tackling with domain shift problem essential for maintaining the robustness and effectiveness of PHM systems. Transfer learning has emerged as a primary solution to domain shift challenges, as it can achieve the adaptation of models from one domain to be effective in another through fine-tuning and domain adaptation techniques. Additionally, robust modeling approaches and the ability to continuously update models with new data are also useful and imperative for mitigating the impact of domain shift.
A clear example of domain shift was observed in the rock drill dataset from the 2022 PHM NA competition. The dataset encompassed training data from five rock drill machines, validation data from another machine, and test data from two additional machines, highlighting potential domain shift issues between training, validation, and test data. Similarly, the 2020 PHM EU dataset demonstrated domain shift, with training data featuring small and large particle sizes, while the test data included medium particle sizes. To address these challenges, various strategies were employed. In the 2022 PHM NA competition, Kim et al. utilized a domain adversarial neural network along with the minimization of maximum mean discrepancy (MMD) to tackle domain discrepancies [31]. Meanwhile, Oh et al. employed the deep CORAL method, calculating coral loss to diminish domain discrepancy effects by aligning the covariance of the six training domains with that of the test domain [53]. For the 2020 PHM EU competition, Tian’s team developed a novel Transfer Ensemble Learning (TEL) framework, facilitating knowledge transfer from the source to the target domain [72]. These approaches illustrate the innovative methodologies being implemented to overcome domain shift problems in PHM, ensuring that models remain accurate and reliable despite changes in their application environments.
@@ 5.2.1 Model Selection: Conventional Machine Learning or Deep Learning?
In PHM data challenge competitions, participants face the critical decision between employing conventional machine learning algorithms or opting for advanced deep learning models. This choice is influenced by various factors: the volume and quality of available data, and the complexity or nature of the research task, whether it be classification or prediction. An analysis of research papers from the past six years reveals insightful trends and preferences in model selection:
The availability of training data plays a significant role in determining the choice of modeling approach. Research teams often choose deep learning for building data-driven models when there is an abundance of training data. Conversely, in scenarios with limited training data, conventional machine learning methods, augmented by feature engineering, are more commonly employed. From Table 3 and Table 5, this trend is evident in the classification challenges of the 2022 PHM NA and 2023 IEEE competitions, where the large datasets provide opportunities to deploy deep learning approaches. However, in the 2022 PHM EU competition, despite sufficient data, the predominance of healthy class data over faulty data led teams to prefer traditional machine learning. In competitions like the 2021 PHM EU and 2023 PHM AP, where training data was around 100-200 samples, traditional machine learning methods were more viable due to deep learning’s requirement for a large volume of data. For RUL prediction tasks, despite smaller datasets (less than 100 samples), the nature of the data being a long time series allows for treating each time point or a series of time points (data window) as a separate training point, effectively enlarging the dataset. Thus, from table 7, we can find that both deep learning and conventional machine learning methods are utilized, except in the 2019 PHM NA competition, where the absence of wave data in part of the test data necessitated a combination of data-driven (conventional machine learning) and physics-based approaches.
@@
In both classification and RUL prediction tasks, Table 4 and Table 7 show that tree-based methods (such as Decision Trees, Random Forest, and XGBoost) and ensemble learning are favored in traditional machine learning. These methods are known for their robustness, ability to handle various data types, and effectiveness in capturing non-linear relationships. Deep learning solutions often favor LSTM-based and CNN-based approaches, with LSTM’s proficiency in handling sequential, time-series data and CNN’s strength in feature extraction from multidimensional datasets. When domain shift is present in the competition dataset, transfer learning and domain adaptation methods are often proposed to adjust the model to new data distributions, leveraging pre-trained models to address limitations in training data. This adaptability ensures the model’s applicability and effectiveness under varied operational conditions, a frequent occurrence in PHM scenarios.
@@ 5.2.2 Machine Learning Model Interpretability
In PHM data challenge competitions, accuracy in prediction and classification is still the primary goal. However, with the advance of deep learning, emphasizing the importance of model interpretability is becoming as crucial as their accuracy in prediction and classification due to its significant impact on safety, operational decision-making, and compliance with industry standards. Model interpretability not only fosters trust and facilitates collaboration between data scientists and domain experts, but it also plays a pivotal role in error analysis and model refinement, which are especially critical in the high-stakes environment of PHM.
Despite the complexity of deep learning models, often referred to as ”black boxes,” various techniques have been developed and applied to enhance their interpretability [66, 67]. Layer-wise Relevance Propagation (LRP), Gradient-weighted Class Activation Mapping in CNNs, and attention mechanisms in sequential models are some of the methods utilized to shed light on model decision-making processes. Additionally, model-agnostic methods like LIME and SHAP have been instrumental in offering insights into model behavior. From a broader prospect, the challenge in PHM is to balance the trade-off between model accuracy and interpretability. Achieving this balance is not just about enhancing model performance, but also about building trust and ensuring compliance with industry standards. Furthermore, in an era increasingly focused on ethical and responsible AI, transparent and interpretable models are key to not only enhancing the technical aspects of AI solutions in PHM but also extending to ensuring their successful integration and acceptance in real-world applications.
@@ 5.2.3 Model Robustness and Generalization
In the realm of PHM, model robustness and model generalization are critical factors that determine the effectiveness of a model in real-world applications. Model robustness is defined as the ability of a model to maintain its performance when facing diverse challenges like noise, outliers, and adversarial examples. Techniques such as data augmentation, where training data can be expanded by creating modified versions of existing data or synthesizing new data, have become commonplace. For instance, in the 2023 IEEE competition, Kreuzer’s team employed methods like additive white Gaussian noise, circular shift, and random amplitude scaling to increase the volume of training data [33]. Moreover, regularization techniques like L1 and L2 regularization have been instrumental in preventing overfitting and enhancing stability, making the model less sensitive to small fluctuations in input data. What’s more, Ensemble methods have been increasingly recognized for their significant contribution to model robustness in PHM. A comprehensive analysis of recent studies shows that out of sixty research papers, nine utilized various ensemble techniques - including ensemble LSTM, decision trees, random forest, XGBoost, soft voting ensemble, transfer ensemble learning, ensemble regression, and ensemble CNN-based approaches - across six different PHM data challenge competitions. This trend underscores the effectiveness of ensemble methods in bolstering model robustness, particularly by reducing the influence of noise and outliers, thereby enhancing the reliability of predictive models in PHM applications. Additionally, adversarial training, which involves training models on both regular and adversarial data, has been recognized for its potential to fortify models against adversarial attacks.
@@
Alongside robustness, model generalization has also been a focal point in PHM, with the goal being to develop models that perform reliably on new, unseen data. One key technique for improving generalization is cross-validation, which involves dividing the training data into several subsets and each time using one of the subsets as the validation data and others as the training data for model validation during the training process. In a notable example from the 2021 PHM NA competition, researchers opted for k-fold repeated random subsampling validation. This approach was chosen as an alternative to standard k-fold cross-validation to address its limitation, wherein the size of the validation set diminishes as the number of folds (k) increases [65]. In addition to cross-validation, transfer learning and domain adaptation methods have been crucial in maintaining model effectiveness and generalization across varied data distributions. These methods are particularly relevant in scenarios characterized by ”domain shift,” where the data from the target domain differs from the source domain. The specific implementations of these methodologies in various competitions have been detailed in the earlier subsubsection titled “Domain Shift”. For deep learning models, specific techniques have been increasingly adopted to improve model generalization and avoid overfitting. These include 1) dropout, where randomly selected neurons are ignored during the training stage; 2) batch normalization, which stabilizes and accelerates training by normalizing layer inputs; 3) hyperparameter optimization, where parameters such as learning rate, number of layers, and batch size are fine-tuned for optimal performance; 4) early stopping, a method where training is halted as soon as the model’s performance on a validation set begins to decline. In summary, these strategies play a critical role in enhancing the robustness and generalizability of machine learning models in the PHM domain.
6 Discussion and Prospects
Our comprehensive analysis of machine learning and deep learning methods across nine open-source industrial datasets has revealed the transformative potential that ML approaches can have on PHM. The nature of the competition tends to prioritize solutions that chiefly enhance accuracy, potentially at the expense of systematic approach, reusability, and methodological inheritance. However, the exploration of methodologies extends outside of the competition. This includes extensive research and analysis of open-source datasets, leading to the development of advanced ML and DL techniques, not just for competitive purposes but for benchmarking and comparative analysis. These efforts have notably augmented the systematization, interpretability, and applicability of ML and DL approaches, thereby expanding their utility beyond the confines of the competition-centric paradigm. Such applications have the power to bridge the gap between academic research and industrial practice, expediting the integration of data-driven machine learning for broader applications across additional industrial datasets.

### v2

@@ 5.1.1 Missing Data
Handling missing data in ML, especially in the context of the PHM is significant, as incomplete data can significantly impact the model performance and prediction accuracy. The strategies to address missing data in PHM have evolved, encompassing a range of techniques from basic deletion to advanced imputation methods. This was evident in the 2021/2022 PHM EU competitions, where missing data posed significant challenges. In the fuse test bench dataset (2021 PHM EU), about 10% of the data was missing, and it was not evenly distributed across variables. Meanwhile, the SPI dataset(2022 PHM EU) had missing information in specific fields. To tackle these issues, various approaches were employed.
A straightforward strategy is the deletion of data points with missing values, such as listwise or pairwise deletion. For example, in the 2022 SPI dataset, instances having missing values in crucial identifiers like ”Panel_ID”, ”Figure_ID”, and ”Component_ID” were eliminated [102]. However, this approach can lead to the loss of valuable information to some degree. On the imputation front, several innovative methods were used in the 2021 PHM EU competition. One team applied the Last Observation Carried Forward (LOCF) method, coupled with backward filling, to address the gaps [22]. Interpolation was another technique used to estimate missing values, ensuring that cases with absent data did not skew the results. Additionally, Ramezani’s team introduced a novel approach, PLS-MV, a partial least squares-based method for imputing missing values [85]. In addition to various imputation methods highlighted in the competitions, there remains scope for exploration in future research such as mean/median/multiple, K-Nearest Neighbors(KNN), regression imputation, maximum likelihood estimation, and ML-based approaches [39]. These methods offer a promising avenue for further enhancing data quality and model reliability in PHM systems.
@@ 5.1.2 Data Imbalance
Data imbalance arises when the distribution of classes in a dataset is uneven. This issue is especially pronounced in PHM due to the rarity of failure events in comparison to data representing normal operations. Data imbalance can significantly undermine the performance of ML models, particularly in classification tasks. When trained on imbalanced data, models may become biased towards the majority class (normal operation) and may not effectively recognize the minority class (failure). This leads to poor performance in predicting failures, which may bring ineffective maintenance planning and unexpected downtimes. In PHM data challenge competitions, this issue is commonly encountered. Fuse test bench dataset from the 2021 PHM EU competition and printed circuit board dataset from the 2022 PHM EU competition exemplify this issue. The former displayed an obvious imbalance with only 3-4 experiments representing each fault class against 70 experiments for the healthy class. Similarly, in the printed circuit board dataset, 95% of the SPI data was classified as healthy, highlighting a significant imbalance. Additionally, the 2018 PHM NA competition also faced this challenge in the Ion Mill Etching (IME) dataset, where the faulty data is much less than normal operation data.
To counteract data imbalance in PHM, various strategies have been adopted. Resampling techniques are commonly used in competitions to adjust the dataset to balance the class distribution, either by oversampling the minority class or downsampling the majority class. Synthetic data generation is another approach, as demonstrated by a team in the 2021 PHM EU competition that used the SMOTE to create synthetic samples of the minority class to balance the dataset [22]. Advanced algorithms also play a crucial role in addressing data imbalance. For instance, Tang’s team in the 2022 PHM EU competition used the FIR to control the imbalance ratio in each mini-batch during neural network training [104]. Liu et al. employed transfer learning and domain adaptation techniques in the 2018 PHM NA competition to facilitate knowledge transfer across different fault modes, addressing the issue of insufficient data in specific faults [66]. Moreover, during the 2018 PHM NA competition, many teams utilized Random Forest, an ensemble learning method known for its proficiency in handling imbalanced data, by constructing a forest of decision trees. These approaches collectively demonstrate the diverse and innovative methods employed to mitigate the challenges posed by data imbalance in the PHM field.
Beyond the techniques showcased in the competitions, new DL approaches are emerging to address this issue, such as the semi-supervised information maximizing generative adversarial network [117], the integration of deep residual networks with auxiliary classifier generative adversarial networks [14], the combination of DL with SMOTE [21], etc. Furthermore, Aguiar et al. introduced a standardized experimental framework to evaluate 24 state-of-the-art data stream algorithms across 515 imbalanced data streams [1]. Going forward, there is significant potential for the exploration of additional DL-based strategies to enhance the handling of data imbalance in the PHM domain.
@@ 5.1.3 Domain Shift
Domain shift refers to the changes in the data distribution between the training phase (source domain) and the real-world application phase (target domain) of the ML models. This phenomenon frequently occurs when models, initially trained on data from a specific set of machines or under certain conditions, are subsequently applied to different machines or varied operating conditions. Such a shift can markedly affect the performance and reliability of ML models in PHM, making tackling with domain shift problem essential for maintaining the robustness and effectiveness of PHM systems. Transfer learning has emerged as a primary solution to domain shift challenges, as it can achieve the adaptation of models from one domain to be effective in another through fine-tuning and domain adaptation techniques. Additionally, robust modeling approaches and the ability to continuously update models with new data are also useful and imperative for mitigating the impact of domain shift.
A clear example of domain shift was observed in the rock drill dataset from the 2022 PHM NA competition. The dataset encompassed training data from five rock drill machines, validation data from another machine, and test data from two additional machines, highlighting potential domain shift issues between training, validation, and test data. Similarly, the 2020 PHM EU dataset demonstrated domain shift, with training data featuring small and large particle sizes, while the test data included medium particle sizes. To address these challenges, various strategies were employed. In the 2022 PHM NA competition, Kim et al. utilized domain adversarial neural networks (DANN) along with the minimization of MMD to tackle domain discrepancies [49]. Meanwhile, Oh et al. employed the deep CORAL method, calculating coral loss to diminish domain discrepancy effects by aligning the covariance of the six training domains with that of the test domain [80]. For the 2020 PHM EU competition, Tian’s team developed a novel TEL framework, facilitating knowledge transfer from the source to the target domain [107]. These approaches illustrate the potential to overcome domain shift problems in PHM, ensuring that models can maintain accurate predictions despite changes in their working or operation conditions.
Table 8: Summary of Methodologies Solving Data-related Issues in PHM Data Challenge Competitions
Data Issue Potential Solution
Missing Data
(1) Listwise or Pairwise Deletion
(2) Imputation (LOCF, PLS-MV)
(3) Other
Data Imbalance
(1) Resampling
(Oversampling or Downsampling)
(2) Synthetic Data Generation
(SMOTE and Variants)
(3) Transfer Learning Techniques
Domain Shift
(1) Transfer Learning Techniques
(2) DANN
@@ 5.2.1 Model Selection: Conventional Machine Learning or Deep Learning?
In PHM data challenge competitions, participants face the critical decision between employing conventional ML algorithms or opting for advanced DL models. This choice is influenced by various factors: the volume and quality of available data, and the complexity or nature of the research task (whether it be classification or prediction). An analysis of research papers from the past six years reveals insightful trends and preferences in model selection:
(1) The volume and quality of training data play a significant role in determining the choice of modeling approach. Research teams often choose DL for building data-driven models when training data are abundant. Conversely, in scenarios with limited training data, conventional ML methods, augmented by feature engineering, are more commonly employed.
This trend is clear when addressing classification problems within PHM competitions. As depicted in Table 2, Table 5, and Figure 2, 2022 PHM NA and 2023 IEEE competitions, because of their large datasets, have facilitated the application of DL techniques. Specifically, the 2023 IEEE competition saw widespread use of deep CNN models, including ensemble-based CNNs [62] and residual-based CNNs [91, 51], showcasing advanced DL’s capability to tackle complex problems. Additionally, 2022 PHM NA competition favored diverse DL approaches like metric learning + pseudo label-based DL [80], CNN + DANN [49], X-Vectors [65], indicating a preference for integrating transfer learning with DL for enhanced performance. However, the scenario differed for the 2022 PHM EU competition, which provided approximately 2000 samples and encountered data imbalance challenges. Given these constraints, many teams leaned towards tree-based methods known for their robustness and ability to address data imbalance, deploying algorithms like XGBoost [32, 102], LightGBM [102], decision trees [75], and random forest [104, 75]. In competitions with significantly fewer data samples, such as the 2021 PHM EU and 2023 PHM AP, where the dataset sizes were around 100-200 samples, the deployment of DL was impractical due to its requirement for a large volume of data. Instead, conventional ML methods proved more effective. For the 2021 PHM EU, tree-based methods were predominantly used [22, 41, 4, 85, 106], while the 2023 PHM AP competition saw a variety of approaches, including similarity-based [73, 47], rule-based [61], KNN/K-means [47, 2], and tree-based methods [61, 2], emphasizing the adaptability of traditional methods to limited data scenarios.
In the realm of RUL prediction tasks, the inherent long-time series nature of the datasets, even when datasets are smaller (less than 100 samples), allows for different approaches to data utilization and data augmentation. By considering each time point or a sequence of time points (data window) as an independent training sample, the effective size of the dataset can be substantially increased. This is evident from Table 7 and Figure 2, which indicate the application of both DL and conventional ML methods across various competitions, with the notable exception of the 2019 PHM NA competition, where only conventional ML methods were employed. As detailed in Figure 2, competitions such as 2018 PHM NA, 2020 PHM EU, and 2021 PHM NA saw widespread adoption of diverse DL techniques, ranging from CNN [113, 97, 24], RNN-based [38, 118, 107] models to transfer learning [66, 107], Autoencoder-based [42], and transformer-based [124, 68] methods, to tackle RUL prediction challenges. Conversely, the 2019 PHM NA competition’s unique challenge, due to the partial absence of wave data in the test dataset, necessitated a hybrid approach that integrated data-driven (conventional ML) and physics-based approaches [50, 121] to address the problem.
@@
(2) The complexity or nature of the research task influences the choice of modeling approach to some degree. Figure 3 shows that tree-based methods (such as Decision Trees, Random Forest, and XGBoost) are favored in conventional ML and DL solutions often favor LSTM-based and CNN-based approaches. A closer examination of Figure 2, Table 4, and Table 7 reveals key insights. The choice between conventional ML and DL for detection and diagnosis (classification problem) hinges largely on dataset size, as discussed before. Absent this factor, for the RUL prediction problem, the majority of teams opted for DL approaches. This is because RNN-based methods (LSTM, GRU, etc) are proficient in handling sequential, time-series data, and CNN-based methods are adept at feature extraction from multidimensional datasets, enhancing the capability to tackle RUL prediction problems efficiently. While conventional ML methods have also shown efficacy, the scalability of DL indicates its superior potential for RUL prediction, given sufficient data. When domain shift is present in the competition dataset, transfer learning and domain adaptation methods are often proposed to adjust the model to new data distributions, and address limitations in training data. This adaptability ensures the model’s applicability and effectiveness under varied operational conditions, a frequent occurrence in PHM scenarios.
(3) When datasets become publicly accessible, there is a clear shift towards the adoption of more sophisticated DL approaches in PHM data challenge competitions over time. Taking the 2018 PHM NA competition as a case study, our review of 12 published papers utilizing this dataset reveals that the methodologies initially favored by the competition teams largely comprised tree-based methods, SVMs, and basic LSTM algorithms [38, 93]. However, as time goes by, there is a discernible trend towards the development of more complex DL algorithms. This includes but is not limited to, GRU [118], knowledge distillation [126], a two-stage deep transfer learning framework utilizing TCN and DANN [66], TCN combined with LSTM [35], attention mechanisms [35, 68], and transformers [124, 68]. This trend underscores a growing need for advancing DL approaches in the PHM domain, to deepen dataset understanding and enhance model performance. The continuous innovation and refinement of DL methods indicate a vibrant future for PHM, where advanced data-driven analyses pave the way for novel discoveries in the field.
(4) A general discussion on the advantages and limitations of CNN, RNN, and tree-based methods in the context of PHM reveals: (1) CNNs excel in extracting features from multidimensional data but struggle with interpretability and require substantial computational resources. (2) RNNs are adept at handling sequential time-series data, benefiting from their ability to remember information over long sequences, useful for understanding long-term dependencies. These approaches are helpful for health state monitoring and RUL prediction, though their complexity and data demands can be challenging. (3) Tree-based methods are valued for their interpretability, feature importance analysis, ability to handle various data types, and effectiveness in capturing non-linear relationships, but they might not achieve the neural networks’ accuracy on large complex datasets.
In summary, selecting an appropriate ML approach should consider the dataset’s specific characteristics, aiming to balance accuracy, interpretability, and computational efficiency to tackle PHM’s diverse challenges. When training data are sufficient, DL methods typically offer superior performance compared to conventional ML approaches.
@@ 5.2.2 Machine Learning Model Interpretability
In PHM data challenge competitions, accuracy in prediction and classification is still the primary goal. However, with the advance of ML and DL, emphasizing the importance of model interpretability is becoming as crucial as their accuracy in prediction and classification due to its significant impact on safety, operational decision-making, and compliance with industry standards. Model interpretability not only fosters trust and facilitates collaboration between data scientists and domain experts, but it also plays a pivotal role in error analysis and model refinement in PHM.
While only a few teams in competitions have used model interpretability methods to explain the output of their models, outside of competitions, various techniques have been developed and applied to enhance ML model interpretability [64, 6]. As the complexity of deep neural network models, often referred to as ”black boxes”, some previous research utilized interpretable methods like Layer-wise Relevance Propagation (LRP), Gradient-weighted Class Activation Mapping in CNNs, and attention mechanisms in sequential models to shed light on model decision-making processes [98]. Additionally, model-agnostic methods like LIME [88] and SHAP [70, 101] have been instrumental in offering insights into model behavior. For instance, Baptista et al. applied the SHAP model to evaluate the outcomes of three different algorithms (Linear Regression, MLP, and Echo State Network) using the Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset (jet engines) [9]. Moreover, Moradi et al. introduced an interpretable artificial neural network designed for the automatic selection and fusion of features to develop optimal health indicators from data gathered through structural health monitoring (SHM) [77]. From a broader prospect, the challenge in PHM is to balance the trade-off between model accuracy and interpretability. Achieving this balance is not just about enhancing model performance, but also about building trust and ensuring compliance with industry standards. Furthermore, in an era increasingly focused on ethical and responsible AI, transparent and interpretable models are key to not only enhancing the technical aspects of AI solutions in PHM but also extending to ensuring their successful integration and acceptance in real-world applications [112].
@@ 5.2.3 Model Robustness and Generalization
In the realm of PHM, model robustness and model generalization are critical factors that determine the effectiveness of a model in real-world applications. Model robustness is defined as the ability of a model to maintain its performance when facing diverse challenges like noise, outliers, and adversarial examples. Techniques such as data augmentation, where training data can be expanded by creating modified versions of existing data or synthesizing new data, have become commonplace. For instance, in the 2023 IEEE competition, Kreuzer’s team employed methods like additive white Gaussian noise, circular shift, and random amplitude scaling to increase the volume of training data [51]. Moreover, regularization techniques like L1 and L2 regularization have been instrumental in preventing overfitting and enhancing stability, making the model less sensitive to small fluctuations in input data. What’s more, Ensemble methods have been increasingly recognized for their significant contribution to model robustness in PHM. A comprehensive analysis of recent studies shows that out of sixty research papers, nine utilized various ensemble techniques - including decision trees, random forest, XGBoost, ensemble LSTM [7], soft voting ensemble [49], transfer ensemble learning [107], ensemble regression [86], and ensemble CNN-based [62] approaches - across six different PHM data challenge competitions. This trend underscores the effectiveness of ensemble methods in bolstering model robustness, particularly by reducing the influence of noise and outliers, thereby enhancing the reliability of predictive models in PHM applications. Additionally, adversarial training, which involves training models on both regular and adversarial data, has been recognized for its potential to fortify models against adversarial attacks [117, 49, 83].
@@
Alongside robustness, model generalization has also been a focal point in PHM, with the goal being to develop models that perform reliably on new, unseen data. One key technique for improving generalization is cross-validation, which involves dividing the training data into several subsets and each time using one of the subsets as the validation data and others as the training data for model validation during the training process. In a notable example from the 2021 PHM NA competition, researchers opted for k-fold repeated random subsampling validation. This approach was chosen as an alternative to standard k-fold cross-validation to address its limitation, wherein the size of the validation set diminishes as the number of folds (k) increases [97]. In addition to cross-validation, transfer learning and domain adaptation methods have been crucial in maintaining model effectiveness and generalization across varied data distributions. These methods are particularly relevant in scenarios characterized by ”domain shift,” where the data from the target domain differs from the source domain. The specific implementations of these methodologies in PHM competitions have been detailed in the earlier subsubsection titled “Domain Shift”. Outside of PHM data challenges, some novel approaches are proposed to deal with model generalization issues. Matthew Russell and Peng Wang adopted a domain adversarial transfer learning method inspired by generative adversarial networks, utilizing a 1D CNN architecture to predict tool wear on unseen domains using NASA’s milling dataset [115]. Ding et al. developed a multi-source domain generalization learning approach (GRU + Transformer) that can effectively learn useful degradation feature representations from various run-to-failure datasets of internal combustion engine journal bearings across different conditions and predict unseen working conditions well [26]. Furthermore, Ding et al. proposed an adversarial out-domain augmentation (AOA) framework for predicting the RUL of bearings under unseen conditions. The effectiveness of this AOA-based RUL prediction was validated using IEEE PHM Challenge 2012 and XJTU-SY run-to-failure datasets, illustrating its robustness in domain generalization for predictive maintenance [27].
5.3 Limitations of Current PHM Competitions and Opportunities
5.3.1 A Lack of Multi-Modal Machine Learning Analysis
Our review of PHM data challenge competitions reveals a reliance on single-modality data in the competitions, such as pressure signals, currents, vibrations, or images, without incorporating multi-modal datasets for fault diagnosis and prognosis. Multi-modal machine learning (MMML) in PHM refers to capturing complementary information from multiple sources (different data types) to achieve a more comprehensive and precise evaluation of PHM tasks [84, 109]. Despite its potential, it is still underexplored in the PHM domain.
Recently, Jiang et al. leveraged two modality data (vibration and current signals) to develop deep belief networks (DBNs) for diagnosing wind turbine gearbox faults [45]. Su et al. proposed an MMML model using parametric specifications, text descriptions, and images of vehicles to predict five vehicle rating scores [100]. Additionally, Fan et al. evaluate several MMML strategies to create a comprehensive PHM system for coolant pumps in commercial heavy-duty vehicles, utilizing data from onboard signals, multi-dimensional histograms, and categorical variables [31]. Wang et al. proposed a novel method for feature fusion in multimodal data (vibration and torque signals), applying it to diagnose the bearings faults [114].
Literature on MMML for PHM tasks shows that most of the methods are focused on classification problems. Therefore, there is still room for further exploration of the application of MMML to prognosis challenges. Moreover, at the heart of MMML for PHM tasks is getting useful multi-modal latent representations using effective data fusion techniques. While simple concatenation methods are widely used for data fusion in MMML, there’s a need for further exploration of more advanced fusion techniques, such as attention-based or transformer-based mechanisms. For future PHM data challenges, the provision of open-source multi-modal datasets would empower researchers to investigate and apply more advanced MMML techniques, potentially leading to significant advancements in the context of PHM.
5.3.2 A Lack of Adopting Systematic Approaches for Effective PHM Systems Construction
Our analysis of ML and DL methods across nine open-source industrial datasets has revealed the transformative potential that ML approaches can have on PHM. Nevertheless, the nature of the competition tends to prioritize solutions that chiefly enhance accuracy, potentially at the expense of systematic approach, reusability, and methodological inheritance. It is therefore vital to pursue systematic methodologies for constructing effective PHM systems that go beyond the competitive framework. This should involve thorough research and analysis of open-source datasets to advance ML and DL strategies, aiming not just for competition success but also for benchmarking and comparative analysis.
Souza et al. devised an ML-based, data-oriented pipeline for constructing a Prognosis and Health Management System (PHMS) focused on RUL prediction, utilizing semi-supervised ML with Autoencoder, XGBoost, and the SHAP method [99]. Professor Enrico Zio highlighted the primary challenges and future directions for implementing condition-based and predictive maintenance in real-world applications [127]. As shown in Figure 5, Hu et al. offered a new perspective on reviewing PHM efforts by proposing a division of the PHM lifecycle into DEsign, DEvelopment, and DEcision ($DE^{3}$) phases, and showcasing the pivotal activities and challenges within these stages [36]. Additionally, Lee et al. introduced a novel SoQ methodology for multi-stage manufacturing processes. It can help to analyze multi-parameter influences on product quality and model inter-process relationships in multi-stage manufacturing systems [58]. Additionally, Figure 6 shows an Industrial Large Knowledge Model framework proposed by Lee’s team. This framework aimed to utilize the power of LLMs and domain-specific knowledge to address the complex challenges in the PHM domain [59].
Figure 5: Interactions of PHM Phases of Design, Development, and Decision ($DE^{3}$) [36]
Moving forward, developing novel, systematic approaches for PHM systems should be encouraged in future PHM data challenge competitions. These efforts can notably augment the systematization and applicability of ML and DL approaches, thereby expanding their utility beyond the confines of the competition-centric paradigm. Such advancements promise to narrow the divide between academic research and industrial practice, facilitating the broader adoption of data-driven ML across various industrial contexts.
6 Prospects